(1 point) Find y as a function of t if 121y + 22y + y = 0, y(0) = 7. y(0) = 9, y =
Answer #1

Given, 121y+22 y+y=0) The characteristic polynomial of the given differential equation is 121m² +22m+1=0. =121m² +22m+1=0 =So, (0– en tele* 0e !) So, 7=ů +c(1-0) Pues - in y=9e il tere 86 9e 11 +- - te 11